OpenStudy (anonymous):

Help!!!! write the slope intercept form of the equation of each line. 3x - y = 4

OpenStudy (anonymous):

Slope-intercept form looks like this:\[\Huge y = mx+b\]It's like all you have to do is isolate y. Do you know how to?

OpenStudy (anonymous):

i don't know how to isolate y :( that's the part i get stuck at

OpenStudy (anonymous):

ok, 3x is with y, and we don't want this, so subtract both sides by 3x to isolate y.

OpenStudy (anonymous):

Now we have -y = 4 - 3x. You know what to do next?

OpenStudy (anonymous):

dont you divide -y to all sides?

OpenStudy (anonymous):

It's isn't good idea because if you do this, you unisolate y.\[\Large -y = 4 - 3x\] \[\Large \frac{-y}{-y} = \frac{4-3x}{-y}\] \[\Large 1= \frac{4-3x}{-y}\]

OpenStudy (anonymous):

We want to isolate y. We have to deal with negative sign. Do you know how to get rid of it?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

-y is same as (-1)( y ) Does this help?

OpenStudy (anonymous):

oh yes. thanks you (: i appreciate it

OpenStudy (anonymous):

Glad I helped.